Rewrite this formula by factoring out the greatest common factor on the right hand side.
A=P+Prt
I came up with P(r+t) but I do not think this is right when you multiply it out to check it does not come out to the original problem.

Answer by bosun1(3)

A=P+Prt
move p to the left side
therefore A-P=PRT.
DIVIDE BOTH SIDE BY P
A-P/P=PRT/P
CROSS OUT P'S
therefore
A=RT [ FINAL ANSWER.